Mr. D.B. Carpenter Flat 8, Broughton House 28/30 Bullar Road Bitterne Park Southampton, S02 4GS

March 8, 1994

Lloyds Bank, Walsall-Aldridge Branch, The Bridge, Walsall, WS1 1LU

Dear Lloyds Bank,

Why is it not technologically possible to produce a payment card with a space large enough to contain the normal human signature? Please can you send me a new blank card. I’m afraid I may need several before I eventually manage to sign one recognisably.

Yours sincerely,